OpenStudy (rockycatsrule):

somone plz help will metdal and fan Using the ratio of perfect squares method, what issquare root of 22 rounded to the nearest hundredth? Do not use your calculator to answer this question. Use the ratio of perfect squares method.

OpenStudy (rockycatsrule):

@mathstudent55 please

OpenStudy (rockycatsrule):

cahek my answer 46.4

OpenStudy (rockycatsrule):

cheak

OpenStudy (juliannegirltomboy):

46.4 i siad already so dont keep posting questions that are the same

OpenStudy (rockycatsrule):

check

OpenStudy (rockycatsrule):

i want to check to see if u r right

OpenStudy (rockycatsrule):

and if u r right then explain how u got your answer and do not look it up !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (juliannegirltomboy):

The ratio of perfect squares method can tell you that sqrt(22) is between 4 and 5, because 4^2=16<22<25=5^2 Estimating using Newton's method will give you an estimate good to the hundredth and more, as follows. Find the closest integer estimate. 4^2=16=22-6, 5^2=25=22+3, so 5 is closer (but too big). We find the estimate by calculating 5-(5^2-22)/(2*5)=4.70 which is almost good to the hundredth. To get more accuracy, we use 4.7 as the estimate: 4.7-(4.7^2-22)/(2*4.7)=4.6904 but then you change them around to get 46.4 the othe numbers don't matter.

OpenStudy (rockycatsrule):

@mathstudent55

OpenStudy (juliannegirltomboy):

i just explained